OpenStudy (anonymous):

what is 10% of 7/8 of160

OpenStudy (anonymous):

14?

OpenStudy (igreen):

10% = 0.10 7/8 = 0.875 0.10 \(\times\) 0.875 = 0.0875 0.0875 \(\times\) 160 = 14

OpenStudy (anonymous):

:D

sammixboo (sammixboo):

Nice job.

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 } = \frac{ x }{ 160 }= 8x=1120 = 140\] \[\frac{ x }{ 140 }\frac{ 10 }{ 100 } 100=1400 = 14\]